# Andreev spin qubit protected by Franck-Condon blockade

### v1

@@ Abstract
Figure 3: Spin-flip rate in response to an external drive at frequency $\omega$ as a function of the coupling $E_{\textrm{so}}$, using Eq. (12). We set $T=0$, $J_{z}/2\pi=0.2$ GHz, $E_{0}/2\pi=1$ GHz, $E_{c}/2\pi=35$ MHz, $\Gamma=50$ MHz. The results illustrate that in the presence of spin-orbit coupling, the spin-flip transition fans out into a set of plasmonic transitions with spacing $\omega_{p}$.
@@
Experimental considerations. Demonstrating the Frack-Condon blockade requires the realization of a strong coupling regime, $\xi_{0}\gg 1$, which is within reach of current experiments. Indeed, measurements of Andreev spins in InAs nanowires routinely observe $E_{0}/2\pi\sim E_{\textrm{so}}/2\pi\sim 1~{}\mathrm{GHz}$ [8, 9, 23, 14]. In that case, to reach strong coupling it is sufficient to construct a circuit with a small charging energy, $E_{C}/2\pi\ll 1~{}\mathrm{GHz}$. Transmon circuits with values as low as $36~{}\mathrm{MHz}$ have been engineered [30]. For these values, $\xi_{0}\approx 3$ can be expected. Potentially, $\xi_{0}$ can be further increased by finding material platforms leading to higher values of $E_{\textrm{so}}$.
@@
Another challenge is the requirement $T\ll\omega_{p}$, but we believe that this condition can also be met in the lab. For the parameters listed above, we expect $\omega_{p}/2\pi\sim 600~{}\mathrm{MHz}$, that is $\sim 30~{}\mathrm{mK}$, above the $\sim 10~{}\mathrm{mK}$ achievable by modern dilution refrigerators. The feasibility of operating low-frequency transmons with small thermal populations was experimentally demonstrated in Ref. [30].
@@
Two final requirements for the operation of the protected Andreev spin qubit concern the feasibility of qubit readout and gates. The spin can be read out dispersively by coupling the circuit to a readout resonator [10, 23]. To achieve such a spin readout, time-reversal symmetry has to be broken during the measurement, for instance by applying a magnetic field perpendicular to spin-orbit field or by threading magnetic flux through a SQUID loop  [^1] See e.g. the Supplemental Information of Ref. [23] for a demontration of Andreev spin readout via a resonator.. The realization of single-qubit gates is complicated by the suppressed overlap between the transmon ground state wave functions for different spins. Fortunately, the Rabi frequency scales with $\xi_{0}$ more mildly, as $e^{-\xi_{0}/2}$, than relaxation, as $e^{-\xi_{0}}$. We leave a detailed analysis of single and two-qubit gates for future investigation.
@@
Conclusions. We have shown that an Andreev spin qubit can be protected against spin relaxation by shunting it with a large capacitance. We provided a set of specific experimental predictions that distinguish the protected regime from an unprotected one. These include an unusual dependence of the spin relaxation rate on the magnetic field as well as the appearance of additional branches detectable in the two-tone spectrum of the system.
@@
We have highlighted the limitations on the protected regime posed by finite temperature of the device. This thermal fragility can be related to the Franck-Condon principle via the sum rule of Eq. (9). The exponential suppression of spin flips due to the disjoint support of the low-lying plasmon states is accompanied by the simultaneous increase of the matrix elements involving the excited plasmon states. At finite temperature, this diminishes the expected returns of being in a protected regime. A similar trade-off could apply to any protected qubit based on a double-well energy landscape [18].
@@
The protected Andreev spin qubit that we propose has noise strongly biased towards pure dephasing. Such a noise bias may be advantageous for quantum error correction [32, 33]. Along these lines, it will be interesting to merge the ideas presented in our manuscript with those of the recent blueprint for Kramers-protected error correction in Andreev spin qubit arrays [34]. Other extensions are the inclusion of a superinductor [35] into the circuit and the study of the Franck-Condon blockade in multi-terminal junctions [36, 37, 38].
Acknowledgments. We thank Valla Fatemi, Lukas Splitthoff, Spencer Diamond, Heekun Nho, Vladislav Kurilovich, Christian K. Andersen, Arno Bargerbos, Marta Pita-Vidal, Jaap Wesdorp, Maximilian Rimbach-Russ, Aleksandr Svetogorov, Leandro Tosi, Leonid Glazman, and Michel Devoret for useful discussions and feedback on the manuscript.
@@
The research of P.D.K., T.C., and C.G.L.B. was sponsored by the Army Research Office (ARO) under grants no. W911NF-22-1- 0053 and W911NF-23-1-0051, by DARPA under grant no. HR0011-24-2-0346, by the U.S. Department of Energy (DoE), Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARO, DARPA, DoE, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Data availability statement. All data generated and code used in this work are available on Zenodo.

### v2

@@ IV.2 Driven excitation of plasmonic branches
Figure 3: Spin-flip rate in response to an external drive at frequency $\omega$ as a function of the coupling $E_{\textrm{so}}$, using Eq. (16). We set $T=0$, $J_{z}/2\pi=0.2$ GHz, $E_{0}/2\pi=1$ GHz, $E_{c}/2\pi=35$ MHz, $\Gamma/2\pi=50$ MHz. The dark lines visible in the figure correspond to distinct spin-flip transitions, each accompanied by the simultaneous excitations of $k$ plasmons, with increasing values $k=0,1,2,3,\dots$ as indicated. The visibility of each transition depends on $E_{\textrm{so}}$ via corresponding Franck-Condon factors $w_{k}$, as in Eq. (16). The results illustrate that in the presence of spin-orbit coupling, the spin-flip transition fans out into a set of plasmonic transitions with spacing $\omega_{p}$.
V Experimental feasibility
@@
Demonstrating the Frack-Condon blockade requires the realization of a strong coupling regime, $\xi_{0}\gg 1$, which is within reach of current experiments. Indeed, measurements of Andreev spins in InAs nanowires routinely observe $E_{0}/2\pi\sim E_{\textrm{so}}/2\pi\sim 1~\mathrm{GHz}$ [8, 9, 26, 14]. In that case, to reach strong coupling it is sufficient to construct a circuit with a small charging energy, $E_{C}/2\pi\ll 1~\mathrm{GHz}$. Transmon circuits with values of $E_{c}$ as low as $36~\mathrm{MHz}$ have been engineered [33]. For these values, $\xi_{0}\approx 3$ can be expected. Potentially, $\xi_{0}$ can be further increased by finding material platforms leading to higher values of $E_{\textrm{so}}$. Fig. 3 shows that for the parameter values quoted in this paragraph the experimental signatures of the protected regime are accessible within the typical bandwidth ($\approx 1-10$ GHz) of a circuit QED experiment.
@@
Another challenge is the requirement $T\ll\omega_{p}$, but we believe that this condition can also be met in the laboratory. For the parameters listed above, we expect $\omega_{p}/2\pi\sim 600~\mathrm{MHz}$, that is $\sim 30~\mathrm{mK}$, above the $\sim 10~\mathrm{mK}$ achievable by modern dilution refrigerators. The feasibility of operating low-frequency transmons with small thermal populations was experimentally demonstrated in Ref. [33].
@@
Two final requirements for the operation of the protected Andreev spin qubit concern the feasibility of qubit readout and gates. The spin can be read out dispersively by coupling the circuit to a readout resonator [10, 26]. To achieve such a spin readout, time-reversal symmetry has to be broken during the measurement, for instance by applying a magnetic field perpendicular to spin-orbit field or by threading magnetic flux through a SQUID loop  [^1] See e.g. the Supplemental Materiak of Ref. [26] for a demontration of Andreev spin readout via a resonator.. The realization of single-qubit gates is complicated by the suppressed overlap between the transmon ground state wave functions for different spins. Fortunately, the Rabi frequency scales with $\xi_{0}$ more mildly, as $e^{-\xi_{0}/2}$, than relaxation, as $e^{-\xi_{0}}$. We leave a detailed analysis of single and two-qubit gates for future investigation.
VI Conclusions
@@
We have shown that an Andreev spin qubit can be protected against spin relaxation by shunting it with a large capacitance. This provides a circuit-based solution, agnostic to the material platform, to the challenge of improving the relaxation time of Andreev spin qubits. We provided a set of specific experimental predictions that distinguish the protected regime from an unprotected one. These include an unusual dependence of the spin relaxation rate on the magnetic field as well as the appearance of additional branches detectable in the two-tone spectrum of the system.
@@
We have highlighted the limitations on the protected regime posed by finite temperature of the device. This thermal fragility can be related to the Franck-Condon principle via the sum rule of Eq. (13). The exponential suppression of spin flips due to the disjoint support of the low-lying plasmon states is accompanied by the simultaneous increase of the matrix elements involving the excited plasmon states. At finite temperature, this diminishes the expected returns of being in a protected regime. A similar trade-off could apply to any protected qubit based on a double-well energy landscape [21, 35].
@@
The protected Andreev spin qubit that we propose has noise strongly biased towards pure dephasing. Such a noise bias may be advantageous for quantum error correction [36, 37, 38]. The strategy of improving relaxation time has been very successful e.g. in cat qubits [39]. This suggests that improving dephasing time, while beneficial, does not need to be the main bottleneck for the development of Andreev spin qubits. Along these lines, it will be interesting to merge the ideas presented in our manuscript with those of the recent blueprint for Kramers-protected error correction in Andreev spin qubit arrays [40]. Other extensions are the inclusion of a superinductor [41] into the circuit [42] and the study of the Franck-Condon blockade in multi-terminal junctions [43, 44, 45].
Acknowledgements.
@@
We thank Valla Fatemi, Lukas Splitthoff, Spencer Diamond, Heekun Nho, Vladislav Kurilovich, Christian K. Andersen, Arno Bargerbos, Marta Pita-Vidal, Jaap Wesdorp, Maximilian Rimbach-Russ, Aleksandr Svetogorov, Leandro Tosi, Leonid Glazman, and Michel Devoret for useful discussions and feedback on the manuscript. The research of P.D.K., T.C., and C.G.L.B. was sponsored by the Army Research Office (ARO) under grants no. W911NF-22-1- 0053 and W911NF-23-1-0051, by DARPA under grant no. HR0011-24-2-0346, by the U.S. Department of Energy (DoE), Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARO, DARPA, DoE, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Data availability
The data that support the findings of this article are openly available [46].